OpenStudy (anonymous):

What is the sum of a 17–term arithmetic sequence where the first term is 6 and the last term is –90? please help?

OpenStudy (anonymous):

Is there a formula for this?

OpenStudy (dumbcow):

didn't i already help you with this one??

OpenStudy (anonymous):

ya but it was wrong :/

OpenStudy (dumbcow):

-714

OpenStudy (anonymous):

Okay! thank you so much. yeah that's a different answer than the one from last time. thank you

OpenStudy (dumbcow):

i gave you the formula Sn = n(A1+An)/2

OpenStudy (dumbcow):

= 17(6-90)/2 = 17(-84)/2 = -714

OpenStudy (anonymous):

Oh sorry. lol I thought you meant 17 was the answer but now i get it. Thank you for giving me the equation too!